# The first phase of mass transfer in low-mass binaries: neither stable nor a common envelope

### v1

@@ 3 Observations
As an observational sample, we use the list of double WD binaries collected by Munday et al. (2024, 2025). We select the binaries that have both masses determined (either because they are double lined, or because they are eclipsing) and for which both measured masses are below to ensure they are likely helium WDs. More massive stars can also form hybrid WDs (with C/O core and helium mantle) with mass below if they lose their envelope very early. But given the narrow range of periods and the slope of the initial mass function (IMF), they should contribute at most a small fraction. The list consists of 18 objects from the literature plus 11 recently discovered binaries discovered in the double-lined double white dwarf survey (DBL, Munday et al. 2024). For the latter only the two masses and the fact that they are in a (relatively) close binary are known, not yet the orbital period. Because we only use the masses this is no problem for our method. We list the parameters of the systems in Table 1, using the updated masses from Munday et al. (2025). For most systems it is clear which of the two WDs is the younger one, given its higher temperature. However, for some systems the two temperatures are not so different and WDs with different masses have different cooling curves. We use the results of Panei et al. (2007) to assess if the formation order could have been different. Since more massive WDs cool slightly slower, we only need to look at systems where the cooler WD is less massive or very similar in mass. System 2, WDJ0054 has two objects with very similar temperatures and mass which could indicate that WD2 was formed first. A similar conclusion holds for systems 13 (WDJ1144), 16 (WDJ141625) and 24 (J2102). In system 12 (WD1101) the two WD have very similar temperature/brightness, so we add this to the list of systems with uncertain formation order. As we will see below, this uncertainty does not influence our conclusions.
@@ 4.1 Reconstruction
As an illustration of the reconstruction process, we show in Fig. 1 the example for WD0455-295 (source id 6 in Table 1). The blue dashed lines show for Solar- (top) and low-metallicity (bottom). At the observed mass of the first formed, currently cooler WD ( ) the range of is indicated by the blue bar (535-735 d). The solid orange lines represent the reconstructed periods , as a function of WD mass, i.e. the period in which the secondary fills its Roche lobe with a core mass equal to the x-coordinate. The upper line is for a secondary mass of 1 at Solar metallicity, the bottom line for 2 star at intermediate metallicity. The reconstructed period range for WD0455-295 is 110-245 d and is shown as the orange bar at the mass of the second formed, currently hotter, WD ( ). The horizontal dotted line shows the expected period at Roche-lobe overflow if the first mass transfer had been stable. It is clear that that is not correct.
We now do the same for all sources, but rather than plotting a figure for each source, we put them all in one plot. In Fig. 2 we show for each source a diamond shape with the range of on the x-axis (equivalent to the blue bar in Fig. 1) and on the y-axis the range of (equivalent to the orange bar in Fig. 1), where we take into account that the metallicities should be the same for both first and second mass transfer. The top right corner of the diamond (with the black dot) is for Solar metallicity and a secondary mass of 1 , which we consider the more likely progenitor range. Lowering the metallicity lowers both periods (top left corner), while increasing lowers only the intermediate period reconstructed from the second phase (bottom right corner). The bottom left corner of the diamond thus is for intermediate metallicity and a 2 secondary, which we consider less likely (since there are fewer more massive stars due to the IMF and most of the stars in the Milky Way have Solar metallicity). We shade the top of the diamonds darker to indicate this higher likelihood.
If all sources had been formed through stable mass transfer, we would expect all the diamonds to intersect with the diagonal . This is clearly not the case, although some sources are consistent with the first phase being stable. In this analysis we did not take the uncertainties in the masses of the WDs into account, which is done in the next section.
@@ 4.2 Forward modelling
We now look at the position in the parameter space where the close double WDs would end up if the first phase of mass transfer was either stable, a common envelope, or governed by the angular momentum balance . For the common-envelope efficiency we use , and for the angular momentum balance we use (see Sect. 4.3). In Fig. 3 we show the resulting distribution plots for the different models, over-plotted with the data points, where we now can take the uncertainties of the data into account. We stress again that our sampling of the progenitor parameter range does not take the relative likelihood like the IMF into account. Similarly, we do not consider selection effects on the parameters of the observed binaries. The question here is simply whether the different models can reach the combination of masses of the observed systems or not.
@@
Regarding the assumption that the first phase was a common envelope (light salmon coloured distribution), we confirm the earlier findings that this is a very poor fit to the data. In these models, the fainter WD is always significantly more massive than the brighter WD, while in the observations this is very rare. Indeed, only SMSS J0338 (id 5), ZTF J0538 (id 7) and ZTF J1749 (id 19) have masses that are compatible with a two common envelope evolution.
@@
Conversely, the model in which the first phase of mass transfer was stable (dark blue distribution) predicts a narrow range of masses, in which the bright WD is somewhat more massive than the faint WD. Note that this region is set by the core mass – radius relation and hardly depends on the mass transfer physics such as how conservative the mass transfer is (see also Zhang et al. 2024). Although about half of the systems have masses compatible with this prediction, it is clear that overall the model is not a good match with the full set of observations, consistent with what we found above. The detailed models of Zhang et al. (2024) find slightly lower mass ratios at low WD mass. Their mass range is shown as the orange area underneath the blue and does not change, or rather supports our conclusion. The parameter range covered by the model using the prescription (green distribution) fits better with the data. One system, SDSS J0634 cannot easily be explained by any of the models as its brighter WD being is much more massive than the fainter one.
@@
We use ConTEST (Stoppa et al. 2023) to do a formal consistency test of the data with respect to the models, despite the fact that we did not take the initial distributions of the progenitor parameters into account. The detailed results are in appendix C and show that the stable mass transfer and double CE model are firmly rejected (p-value 0.0), while the model using the prescription is formally consistent with a p-value of 0.13, if we exclude the data points that are consistent with two common envelopes (ids 5, 7 and 19). Note that we had to remove all WD2 masses below in the models since there is no support for these values in the data, which may be due to selection effects.
@@ 4.3 Reconstructing values
Following up on the previous finding that most systems can be explained by a model using the prescription we reconstruct the values of for each of the observed systems, taking the observed masses into account, sampling the uncertainties assuming a normal distribution and flat distribution of initial masses between 1 and 2 . The results are plotted in Fig. 4, where we show a kernel density estimate of 120 random selections from the parameter distribution (80 for Solar and 40 for intermediate metallicity). It confirms that all systems apart fron system 8 are compatible with a value of between 1.5 and 1.75. We note that the simulation samples the full uncertainty, so contains systems outside the 1 region. That is the reason why e.g. source 5 falls far from the green area in Fig. 3, but still has solutions with .
The possible solutions of assuming the first phase of mass transfer also was a common envelope are shown in Appendix B. For 13 systems there are no solutions with positive values, while for another 8 there are only solutions with . This confirms the conclusion that two common-envelope phases are a poor fit to the observations.
@@ 5 Discussion and conclusions
In our analysis, we have made a number of assumptions that could affect the results. First, we use an analytic core mass – radius relation, rather than performing detailed stellar evolution calculations. Previous work has shown that the final outcome of stable mass transfer typically follows the core mass – radius relation tightly (e.g. Rappaport et al. 1995; Tauris & Savonije 1999; Lin et al. 2011). The recent detailed study of Zhang et al. (2024) is exactly targeted at the evolution we study. They do find that the mass ratio of the final double helium WD is somewhat lower at low WD mass compared to their semi-analytic model ( instead of ). They conclude that this is due to small additional hydrogen mass on top of the core (making the first WD more massive) as well as small deviations of the core mass – radius relation for stars that have started mass transfer before or at the base of the red giant branch. Interestingly, our semi-analytical model, using a slightly different core mass – radius relation, gives mass ratios of at low WD mass, so in between their semi-analytic and detailed models. In Fig. 3 we indicate the region of the detailed models of Zhang et al. (2024) in orange showing the small offset, but supporting our conclusion.
@@
Gao & Li (2023) investigate the effect of stellar wind and find that normal winds do not change the core mass – period relation significantly, but strong tidally enhanced stellar winds can lead to significantly wider periods, as the donor can lose Roche-lobe contact. In principle this could widen the parameter range from stable mass transfer, but only towards more unequal masses, not to systems with equal masses or reversed mass ratios.
@@
More generally, we assume that the period of the binary does not change between the end of the first phase of mass transfer and the onset of the second phase of mass transfer. Apart from (strong) winds, this could also be influenced by tides, if there is significant transfer of orbital angular momentum to the spin of the secondary. This would shrink the orbit and bring the mass of the second formed WD closer to the mass of the first formed WD, but would preserve the tight relation.
@@
Additionally, we have assumed that all observed double WD systems with mass below originate from mass transfer of stars with initial masses between 1 and 2 , that start mass transfer after the main sequence (case B). The relations we derive do not hold for more massive stars (that do not have degenerate helium cores) and for mass transfer that starts on the main sequence. A full population model will thus add different formation channels for double WD that add systems to the mass range we consider and will likely populate other regions. The IMF and initial period distribution suggest that the majority of systems follow the evolution outlined here. However, there seems to be a cluster of systems with relatively high masses (total mass around 0.9), perhaps indicating a contribution from more massive progenitors. Furthermore, we have used only two metallicity models. Using a broader range of metallicities would broaden the range of periods that are possible after stable mass transfer. However, our metallicities cover the majority of the star formation that has happened in our Milky Way, so we do not believe that this will change the conclusions of our paper. Finally, we assume that the mass of the core does not grow during the second phase of mass transfer, which we believe is well justified by the expected short duration of the common-envelope phase.
@@
Our results seem at odds with the conclusions of Ge et al. (2024); Zhang et al. (2024) that the formation of double WD and sdB + WD binaries points at a first mass transfer that is stable. They also use the core mass – radius relation to infer the periods before the second phase (which is assumed to be a common envelope) and they select the progenitor from a set of models that follow the core mass – radius relation. Zhang et al. (2024) compare their resulting mass ratio’s qualitatively with those measured in double WD and for our systems 10 and 11 do a more detailed study (mainly to determine the efficiency of the final common envelope). These two indeed are consistent with stable mass transfer. To study the sdB + WD binaries we plot in Fig. 5 the masses of the sdB and WD in the systems, mainly taken from Schaffenroth et al. (2022, 2023). Ge et al. (2024) assume the two objects with low sdB mass are formed from stars more massive than 2 that have non-degenerate cores and the others could indeed be consistent with the assumption that the first phase was stable, but sometimes there is a lack of agreement (e.g. they find periods that differ from the predicted period by up to 60%). In light of our conclusion above that the first phase of mass transfer cannot always have been stable for double WD that follow very much the same evolution, we believe the sdB + WD binaries likely also did not all go through a first phase that is stable. Indeed, the distribution of data points in Fig. 5 is similar to that in Fig. 3. Zhang et al. (2021) used a very similar method to ours to construct the final helium WD mass - period relation for helium WD + sdB binaries. They concluded that stable mass transfer is in broad agreement with the observations. However, they make the comparison more complicated by including the final common-envelope phase and their model proves not to fit the data very well.
@@
Li et al. (2023) simulate the Galactic population of double WD using stability criteria that favour stable mass transfer and find that the results are a smaller population that fits the observational constraints better. Surprisingly, the resulting mass ratio distribution of the double WD (their Figure 7) has a strong peak at equal masses, not at , in contrast to our results. This may be related to the treatment of stable mass transfer.
@@
Di Stefano et al. (2023) suggest an approach to the common envelope that is also based on angular momentum, but with a more physical interpretation. The assumption is that the angular momentum that is taken with the material lost from the binary is the sum of what is provided by the core of the giant and what is provided by the companion. The model depends on a parameter that multiplies the specific the angular momentum that is transferred to the envelope (for an assumed value of the dimensionality parameter ). Based on known post common-envelope systems, Di Stefano et al. (2023) find that is large and a function of the ratio of the total mass and the envelope mass with systems with larger ratio having larger . We investigate the results if the first phase of mass transfer could be described by their fit to and show the resulting mass distribution in Fig. 6 as the light pink distribution. Since this relation was fitted to systems that underwent clear shrinkage of their orbit, it is perhaps not surprising that the results with such high values look very much like the case where the first phase of mass transfer was described by the standard common envelope equations. We do find that if the first phase of mass transfer was described by this formalism but with a constant value , the resulting masses (dark cyan distribution in Fig. 6) are similar to those found for the formalism and even somewhat more extended, potentially covering more of the observed systems.
@@
In summary, we have used two methods to investigate if the first phase of mass transfer leading to double helium WD binaries was stable and conclude that this generally is not the case. Where stable mass transfer is expected to lead to a well-defined period after the first mass transfer and a strong correlation between the masses of the final two WDs, the observations suggest a very broad range of periods after the first phase of mass transfer. Despite its lack of clear physical interpretation, the prescription still gives the best match of the models to the observations.
@@
The mismatch between the periods after the first mass transfer and models assuming stable mass transfer may very well be related to the finding that many of the observed binaries that are in the phase (just) after the first mass transfer have eccentric orbits, while the previous mass transfer is expected to lead to circular orbits for all models we consider here (e.g. Bonačić Marinović et al. 2008; Dewberry & Wu 2025). Several solutions have been investigated (e.g. Bonačić Marinović et al. 2008; Dermine et al. 2013; Vos et al. 2015; Saladino & Pols 2019; Escorza et al. 2020; Oomen et al. 2020), many of which have problems and several rely on the donor having been an AGB star undergoing strong stellar-wind mass loss. However, the fact that also descendants of RGB donors can have eccentric orbits and as we show here do not seem to follow stable mass transfer, suggests the problem is not unique to AGB donors. The solution to that conundrum perhaps can also explain the masses of the observed double WD population.

### v2

@@ 3 Observations
As an observational sample, we used the list of double WD binaries collected by Munday et al. (2024, 2025). We selected the binaries that have had both masses determined (either because they are double lined or because they are eclipsing) and for which both measured masses are below to ensure they are likely helium WDs. More massive stars can also form hybrid WDs (with C/O core and helium mantle) with masses below if they lose their envelope very early. However, given the narrow range of periods and the slope of the initial mass function (IMF), they should contribute at most a small fraction. The list consists of 18 objects from the literature and 11 recently discovered binaries from the double-lined double-WD survey (DBL, Munday et al. 2024). For the latter, only the two masses and the fact that they are in a (relatively) close binary are known; the orbital period is not yet known. Because we only used the masses, this is not a problem for our method. We listed the parameters of the systems in Table 1, using the updated masses from Munday et al. (2025). For most systems, it is clear which of the two WDs is the younger one, given its higher temperature. However, for some systems the two temperatures are not so different, and WDs with different masses have different cooling curves. We used the results of Panei et al. (2007) to assess if the formation order could have been different. Since more massive WDs cool slightly more slowly, we only needed to look at systems where the cooler WD is less massive or very similar in mass. System 2, WDJ0054, has two objects with very similar temperatures and mass, which could indicate that WD2 was formed first. A similar conclusion holds for systems 13 (WDJ1144), 16 (WDJ141625), and 24 (J2102). In system 12 (WD1101), the two WDs have very similar temperature/brightness, so we added this to the list of systems with uncertain formation orders. As we see below, this uncertainty does not influence our conclusions.
@@ 4.1 Reconstruction
As an illustration of the reconstruction process, we show, in Fig. 1, the example for WD0455-295 (source ID 6 in Table 1). The dashed blue lines show for solar- (top) and low-metallicity (bottom). At the observed mass of the first formed, currently cooler WD ( ), the range of is indicated by the blue bar (535-735 d). The solid orange lines represent the reconstructed periods, , as a function of WD mass (i.e. the period in which the secondary fills its Roche lobe with a core mass equal to the x-coordinate). The upper line is for a secondary mass of 1 at solar metallicity, and the bottom line is for 2 star at intermediate metallicity. The reconstructed period range for WD0455-295 is 110-245 d and is shown as the orange bar at the mass of the second formed, currently hotter WD ( ). The horizontal dotted line shows the expected period at Roche-lobe overflow if the first mass transfer had been stable. It is clear that that is not correct.
We then repeated the procedure for all sources, but rather than plotting a figure for each source, we put them all in one plot. In Fig. 2, for each source we show a diamond shape with the range of on the x-axis (equivalent to the blue bar in Fig. 1) and the range of on the y-axis (equivalent to the orange bar in Fig. 1), where we take into account that the metallicities should be the same for both first and second mass transfer. The top right corner of the diamond (with the black dot) is for Solar metallicity and a secondary mass of 1 , which we consider the more likely progenitor range. Lowering the metallicity lowers both periods (top left corner), while increasing lowers only the intermediate period reconstructed from the second phase (bottom right corner). The bottom left corner of the diamond is thus for intermediate metallicity and a 2 secondary, which we consider less likely (since there are fewer more massive stars due to the IMF and most of the stars in the Milky Way have Solar metallicity). We shade the top of the diamonds darker to indicate this higher likelihood.
If all sources had been formed through stable mass transfer, we would expect all the diamonds to intersect with the diagonal . This is clearly not the case, although some sources are consistent with the first phase being stable. In this analysis, we did not take the uncertainties in the masses of the WDs into account; this is done in the next section.
@@ 4.2 Forward modelling
We now look at the position in the parameter space where the close double WDs would end up if the first phase of mass transfer were stable, a common envelope, or governed by the angular-momentum balance . For the common-envelope efficiency, we used ; and for the angular-momentum balance, we used (see Sect. 4.3). In Fig. 3, we show the resulting distribution plots for the different models, overplotted with the data points, where we can now take the uncertainties of the data into account. We stress again that our sampling of the progenitor parameter range does not take a relative likelihood such as the IMF into account. Similarly, we did not consider selection effects on the parameters of the observed binaries. The question here is simply whether the different models can reach the combination of masses of the observed systems or not.
@@
Regarding the assumption that the first phase was a common envelope (light salmon coloured distribution), we confirm the earlier findings that this is a very poor fit to the data. In these models, the fainter WD is always significantly more massive than the brighter WD, while in the observations this is very rare. Indeed, only SMSS J0338 (ID 5), ZTF J0538 (ID 7), and ZTF J1749 (ID 19) have masses that are compatible with a two-common-envelope evolution.
@@
Conversely, the model in which the first phase of mass transfer was stable (dark blue distribution) predicts a narrow range of masses in which the bright WD is somewhat more massive than the faint WD. We note that this region is set by the core mass – radius relation and hardly depends on the mass transfer physics such as how conservative the mass transfer is (see also Zhang et al. 2024). Although about half of the systems have masses compatible with this prediction, it is clear that overall the model is not a good match with the full set of observations; this is consistent with what we found above. The detailed models of Zhang et al. (2024) find slightly lower mass ratios at low WD masses. Their mass range is shown as the orange area underneath the blue one and does not change; rather, it supports our conclusion. The parameter range covered by the model using the prescription (green distribution) fits better with the data. One system, SDSS J0634, cannot easily be explained by any of the models as its brighter WD is much more massive than the fainter one.
@@
We used ConTEST (Stoppa et al. 2023) to carry out a formal consistency test of the data with respect to the models, despite the fact that we did not take the initial distributions of the progenitor parameters into account. The detailed results are given in Appendix C and show that the stable mass transfer and double CE model are firmly rejected (p-value 0.0), while the model using the prescription is formally consistent with a p value of 0.13 if we exclude the data points that are consistent with two common envelopes (IDs 5, 7, and 19). We note that we had to remove all WD2 masses below in the models since there is no support for these values in the data, which may be due to selection effects.
To assess the impact of a possible reversal of the masses in the systems in which the two WDs have very similar temperatures and ages, we plot five systems (systems 2, 12, 13, 16, and 24 in Table 1), with the masses swapped in blue in Fig. 3. The conclusions do not change. In fact, the situation worsens.
@@ 4.3 Reconstructing values
Following up on the previous finding that most systems can be explained by a model using the prescription, we reconstructed the values of for each of the observed systems, taking the observed masses into account, sampling the uncertainties assuming a normal distribution and flat distribution of initial masses between 1 and 2 . The results are plotted in Fig. 4, where we show a kernel density estimate of 120 random selections from the parameter distribution (80 for solar and 40 for intermediate metallicity). It confirms that all systems apart from system 8 are compatible with a value of between 1.5 and 1.75. We note that the simulation samples the full uncertainty, so it contains systems outside the 1 region. That is why source 5, for example, falls far from the green area in Fig. 3, but still has solutions with .
The possible solutions of assuming the first phase of mass transfer was also a common envelope are shown in Appendix B. For 13 systems, there are no solutions with positive values, while for another eight there are only solutions with . This confirms the conclusion that two common-envelope phases are a poor fit to the observations.
@@ 5 Discussion and conclusions
In our analysis, we made a number of assumptions that could affect the results. First, we used an analytic core mass – radius relation, rather than performing detailed stellar evolution calculations. Previous work has shown that the final outcome of stable mass transfer typically follows the core mass – radius relation tightly (e.g. Rappaport et al. 1995; Tauris and Savonije 1999; Lin et al. 2011). The recent detailed study of Zhang et al. (2024) is exactly targeted at the evolution we studied. They do find that the mass ratio of the final double-helium WD is somewhat lower at a low WD mass compared to their semi-analytic model ( instead of ). They conclude that this is due to a small amount of additional hydrogen mass on top of the core (making the first WD more massive) as well as small deviations of the core mass – radius relation for stars that started mass transfer before or at the base of the red giant branch. Interestingly, our semi-analytical model, using a slightly different core mass – radius relation, gives mass ratios of at low WD masses; that is, between their semi-analytic and detailed models. In Fig. 3, we indicate the region of the detailed models of Zhang et al. (2024) in orange, showing the small offset but supporting our conclusion.
@@
Gao and Li (2023) investigated the effect of stellar wind and found that normal winds do not change the core mass – period relation significantly, but strong, tidally enhanced stellar winds can lead to significantly wider periods, as the donor can lose Roche-lobe contact. In principle, this could widen the parameter range from stable mass transfer, but only towards more unequal masses and not to systems with equal masses or reversed mass ratios.
@@
More generally, we assumed that the period of the binary does not change between the end of the first phase of mass transfer and the onset of the second phase of mass transfer. Apart from (strong) winds, this could also be influenced by tides if there is significant transfer of orbital angular momentum to the spin of the secondary. This would shrink the orbit and bring the mass of the second formed WD closer to the mass of the first formed one, but it would preserve the tight relation.
@@
Additionally, we assumed that all observed double-WD systems with masses below originate from mass transfer of stars with initial masses between 1 and 2 that begin mass transfer after the main sequence (case B). The relations we derive do not hold for more massive stars (that do not have degenerate helium cores) and for mass transfer that starts on the main sequence. A full population model will thus add different formation channels for double WD that add systems to the mass range we considered and will likely populate other regions. The IMF and initial period distribution suggest that the majority of systems follow the evolution outlined here. However, there seems to be a cluster of systems with relatively high masses (total mass around 0.9), perhaps indicating a contribution from more massive progenitors. Furthermore, we only used two metallicity models. Using a broader range of metallicities would widen the range of periods that are possible after stable mass transfer. However, our metallicities cover the majority of the star formation that has happened in our Milky Way, so we do not believe that this will change the conclusions of our work. Finally, we assumed that the mass of the core does not grow during the second phase of mass transfer, which we believe is well justified by the expected short duration of the common-envelope phase.
@@
Our results seem at odds with the conclusions of Ge et al. (2024) and Zhang et al. (2024) that the formation of double WD and sdB + WD binaries points at a stable initial mass transfer. They also used the core mass – radius relation to infer the periods before the second phase (which is assumed to be a common envelope) and selected the progenitor from a set of models that follow the core mass – radius relation. Zhang et al. (2024) compared their resulting mass ratios qualitatively with those measured in double WD, and for our systems 10 and 11 they carried out a more detailed study (mainly to determine the efficiency of the final common envelope). These two are indeed consistent with stable mass transfer. To study the sdB + WD binaries, we plot the masses of the sdB and WD in the systems, mainly taken from Schaffenroth et al. (2022, 2023), in Fig. 5. Ge et al. (2024) assumed the two objects with low sdB mass are formed from stars more massive than 2 that have non-degenerate cores and that the others could indeed be consistent with the assumption that the first phase was stable, but sometimes there is a lack of agreement (e.g. they find periods that differ from the predicted period by up to 60%). In light of our conclusion above that the first phase of mass transfer cannot always have been stable for double WDs that very much follow the same evolution, we believe the sdB + WD binaries likely also did not all go through a stable first phase. Indeed, the distribution of data points in Fig. 5 is similar to that in Fig. 3. Zhang et al. (2021) used a very similar method to ours to construct the final helium WD mass - period relation for helium WD + sdB binaries. They concluded that stable mass transfer is in broad agreement with the observations. However, they made the comparison more complicated by including the final common-envelope phase, and their model proves not to fit the data very well.
@@
Li et al. (2023) simulated the Galactic population of double WDs using stability criteria that favour stable mass transfer and found that the results provide a smaller population that fits the observational constraints better. Surprisingly, the resulting mass-ratio distribution of the double WD (their Figure 7) has a strong peak at equal masses, but not at , in contrast to our results. This may be related to the treatment of stable mass transfer.
@@
Di Stefano et al. (2023) suggested an approach to the common envelope that is also based on angular momentum, but with a more physical interpretation. The assumption is that the angular momentum that is taken with the material lost from the binary is the sum of what is provided by the core of the giant and what is provided by the companion. The model depends on a parameter that multiplies the specific angular momentum that is transferred to the envelope (for an assumed value of the dimensionality parameter ). Based on known post-common-envelope systems, Di Stefano et al. (2023) found that is large and a function of the ratio of the total mass and envelope mass, with systems with larger ratios having larger . We investigated the results if the first phase of mass transfer could be described by their fit to and show the resulting mass distribution in Fig. 6 as the light pink distribution. Since this relation was fitted to systems that underwent clear shrinkage of their orbit, it is perhaps not surprising that the results with such high values look very much like the case where the first phase of mass transfer was described by the standard common-envelope equations. We do find that if the first phase of mass transfer was described by this formalism but with a constant value , the resulting masses (dark cyan distribution in Fig. 6) are similar to those found for the formalism and even somewhat more extended, potentially covering more of the observed systems.
@@
In summary, we used two methods to investigate if the first phase of mass transfer leading to double-helium WD binaries was stable and conclude that this is not generally the case. Where stable mass transfer is expected to lead to a well-defined period after the first mass transfer and a strong correlation between the masses of the final two WDs, the observations suggest a very broad range of periods after the first phase of mass transfer. Despite its lack of a clear physical interpretation, the prescription still gives the best match of the models to the observations.
@@
The mismatch between the periods after the first mass transfer and models assuming stable mass transfer may very well be related to the finding that many of the observed binaries that are in the phase (just) after the first mass transfer have eccentric orbits, while the previous mass transfer is expected to lead to circular orbits for all models we considered here (e.g. Bonačić Marinović et al. 2008; Dewberry and Wu 2025). Several solutions have been investigated (e.g. Bonačić Marinović et al. 2008; Dermine et al. 2013; Vos et al. 2015; Saladino and Pols 2019; Escorza et al. 2020; Oomen et al. 2020), many of which have problems and several rely on the donor having been an AGB star undergoing strong stellar-wind mass loss. However, the fact that descendants of RGB donors can also have eccentric orbits and, as we show here, do not seem to follow stable mass transfer, suggests that the problem is not unique to AGB donors. The solution to that conundrum can perhaps also explain the masses of the observed double-WD population.